Three point charges of +2 uC, -3uC and -3 uC are kept at the vertices A, B and C respectively, of an equilateral triangle of side 20cm. What should be the sign and magnitude of the charge (q) to be placed at the midpoint (M) of side BC so that the charge at A remains in equilibrium?

Grade:11

1 Answers

Khimraj
5 years ago
sign of q must be positive.
magnitude
[K*2*10-6*3*10-6/(0.2)2]*cos30 + [K*2*10-6*3*10-6/(0.2)2]*cos30 = [K*2*10-6*q/0.03]
2*3*10-6*cos30/0.04 = q/0.03
q = 18*cos30/4 uC
q = 3.9 uC
Hope it clears............................